OpenStudy (anonymous):

Nikki makes $9.50 an hour working at Current Electronics. She plans to buy a DVD recorder that costs $269.60. Write and solve an inequality describing how many hours h Nikki will have to work to be able to buy the DVD recorder.

OpenStudy (anonymous):

@johnweldon1993 here i will right it down here

OpenStudy (anonymous):

269.60 – h ≥ 9.50; 28 hours 269.60 + h ≥ 9.50; 31 hours 269.60h ≥ 9.50; 30 hours 9.50h ≥ 269.60

OpenStudy (anonymous):

\[x+19\le-5\] @johnweldon1993

OpenStudy (johnweldon1993):

Ahh see, well the one you sent had -15 not -5 lol well same thing...subtract 19 from both sides.... \[\large x\le ?\]

OpenStudy (anonymous):

14

OpenStudy (johnweldon1993):

Not quite \[\large -5 - 19 = ?\]

OpenStudy (anonymous):

-24

OpenStudy (anonymous):

\[-0.3n \le5.4\]@Johnyoungman224

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@johnweldon1993